# Project page: https://dai-wenxun.github.io/MotionLCM-page

### v2

@@ Abstract
This work introduces MotionLCM, extending controllable motion generation to a real-time level. Existing methods for spatial-temporal control in text-conditioned motion generation suffer from significant runtime inefficiency. To address this issue, we first propose the motion latent consistency model (MotionLCM) for motion generation, building upon the latent diffusion model [9]. By adopting one-step (or few-step) inference, we further improve the runtime efficiency of the motion latent diffusion model for motion generation. To ensure effective controllability, we incorporate a motion ControlNet within the latent space of MotionLCM and enable explicit control signals (e.g., initial poses) in the vanilla motion space to control the generation process directly, similar to controlling other latent-free diffusion models [73, 29] for motion generation. By employing these techniques, our approach can generate human motions with text and control signals in real-time. Experimental results demonstrate the remarkable generation and controlling capabilities of MotionLCM while maintaining real-time runtime efficiency.
@@ 1 Introduction
Here, in MotionLCM, we are facing another challenge on how to control motions with spatial-temporal signals (e.g., initial poses) in the latent space. Previous methods [56, 73] model human motions in the vanilla motion space and can manipulate the motion directly in the denoising process. However, for our latent-diffusion-based MotionLCM, it is non-trivial to feed the control signals into the latent space. This is because the latent has no explicit motion semantics, which cannot be manipulated directly by the control signals. Inspired by the notable success of [82] in controllable image generation [55], we introduce a motion ControlNet to control motion generation in the latent space. However, the naïve motion ControlNet is not sufficient to provide supervision for the control signals. The main reason is the lack of explicit supervision in the motion space. Therefore, during the training phase, we decode the predicted latent through the frozen VAE [30] decoder into the vanilla motion space to provide explicit control supervision on the generated motion. Thanks to the powerful one-step inference capability of MotionLCM, the latent generated by MotionLCM can significantly facilitate control supervision both in the latent space and motion space for training the motion ControlNet compared to MLD [9].
With our key designs, our proposed MotionLCM successfully enjoys the balance between the generation quality and efficiency in controllable motion generation. Before delivering into detail, we sum up our core contributions as follows.
@@ 2 Related Work
2.1 Human Motion Generation
Generating human motions can be divided into three main fashions according to inputs: motion synthesis (1) without any condition [77, 84, 54, 65], (2) with some given multi-modal conditions, such as action labels [19, 48, 6, 75, 31, 12], textual description [13, 7, 1, 38, 37, 2, 5, 14, 17, 64, 49, 18, 83, 3, 65, 70, 9, 11, 78, 81, 56, 29, 50, 27, 72, 43, 68, 15, 39, 86, 51, 4, 69, 24, 10, 7, 85], audio or music [35, 32, 59, 66, 34, 33], (3) with user-defined trajectories [56, 29, 73, 71, 68, 23, 22, 58, 36]. To generate diverse, natural, and high-quality human motions, many generative models have been explored by [2, 38, 49, 79, 80]. Recently, diffusion-based models significantly improved the motion generation performance and diversity [83, 65, 9, 11, 8, 76, 16] with stable training. Specifically, MotionDiffuse [83] represents the first text-based motion diffusion model that provides fine-grained instructions on body parts and achieves arbitrary-length motion synthesis with time-varied text prompts. MDM [65] introduces a motion diffusion model that operates on raw motion data, enabling both high-quality generation and generic conditioning that together comprise a good baseline for new motion generation tasks. Based on MDM [65], OmniControl [73] integrates flexible spatial-temporal control signals across different joints by combining analytic spatial guidance and realism guidance into the diffusion model, ensuring that the generated motion closely conforms to the input control signals. The work most relevant to ours is MLD [9], which introduces a motion latent-based diffusion model to enhance generation quality and reduce computational resource requirements. The key idea is training a VAE [30] for motion embedding, followed by implementing latent diffusion [55] within the learned latent space. However, these diffusion fashions inevitably require considerable sampling steps for motion synthesis during inference, even with some sampling acceleration methods [61]. Thus, we propose MotionLCM, which not only guarantees high-quality controllable motion generation but also achieves real-time runtime efficiency.
@@ 3.3 Controllable Motion Generation in Latent Space
$\mathbf{\Theta}^{a},\mathbf{\Theta}^{b}=\underset{\mathbf{\Theta}^{a},\mathbf{\Theta}^{b}}{\text{argmin}}({\cal L_{\text{recon}}}+\lambda{\cal L_{\text{{control}}}}),$ (10)
where $\lambda$ is the weight to balance the two losses. This design enables explicit control signals to directly influence the generation process, similar to controlling other latent-free diffusion models for motion generation [73, 29]. Extensive experiments demonstrate the introduced supervision is very helpful in improving the motion control performance, which will be introduced in the following section.
@@ 4.4 Ablation Studies
Table 5: Ablation study on different control ratios $\tau$ and number of control joints $K$. We report the results of (1, 2, 4)-step inference.“” denotes the default training setting.

### v3

@@ Abstract
This work introduces MotionLCM, extending controllable motion generation to a real-time level. Existing methods for spatial-temporal control in text-conditioned motion generation suffer from significant runtime inefficiency. To address this issue, we first propose the motion latent consistency model (MotionLCM) for motion generation, building on the motion latent diffusion model. By adopting one-step (or few-step) inference, we further improve the runtime efficiency of the motion latent diffusion model for motion generation. To ensure effective controllability, we incorporate a motion ControlNet within the latent space of MotionLCM and enable explicit control signals (i.e., initial motions) in the vanilla motion space to further provide supervision for the training process. By employing these techniques, our approach can generate human motions with text and control signals in real-time. Experimental results demonstrate the remarkable generation and controlling capabilities of MotionLCM while maintaining real-time runtime efficiency.
@@ 1 Introduction
Here, in MotionLCM, we are facing another challenge on how to control motions with spatial-temporal signals (i.e., initial motions) in the latent space. Previous methods [56, 73] model human motions in the vanilla motion space and can manipulate the motion directly in the denoising process. However, for our latent-diffusion-based MotionLCM, it is non-trivial to feed the control signals into the latent space. This is because the latent has no explicit motion semantics, which cannot be manipulated directly by the control signals. Inspired by the notable success of [82] in controllable image generation [55], we introduce a motion ControlNet to control motion generation in the latent space. However, the naïve motion ControlNet is not sufficient to provide supervision for the control signals. The main reason is the lack of explicit supervision in the motion space. Therefore, during the training phase, we decode the predicted latent through the frozen VAE [30] decoder into the vanilla motion space to provide explicit control supervision on the generated motion. Thanks to the powerful one-step inference capability of MotionLCM, the latent generated by MotionLCM can significantly facilitate control supervision both in the latent space and motion space for training the motion ControlNet compared to MLD [9].
With our key designs, our proposed MotionLCM successfully enjoys the balance between generation quality and efficiency in controllable motion generation. Before delivering into detail, we sum up our core contributions as follows.
@@ 2 Related Work
Generating human motions can be divided into three main fashions according to inputs: motion synthesis (1) without any condition [77, 84, 54, 65], (2) with some given multi-modal conditions, such as action labels [19, 48, 6, 75, 31, 12], textual description [1, 38, 37, 2, 5, 14, 17, 64, 49, 18, 83, 3, 65, 70, 9, 11, 78, 81, 56, 29, 50, 27, 72, 43, 68, 15, 39, 86, 51, 4, 69, 24, 13, 85, 10, 7], audio or music [35, 32, 59, 66, 34, 33], (3) with user-defined trajectories [56, 29, 73, 71, 68, 23, 22, 58, 36]. To generate diverse, natural, and high-quality human motions, many generative models have been explored by [2, 38, 49, 79, 80]. Recently, diffusion-based models significantly improved the motion generation performance and diversity [83, 65, 9, 11, 8, 76, 16] with stable training. Specifically, MotionDiffuse [83] represents the first text-based motion diffusion model that provides fine-grained instructions on body parts and achieves arbitrary-length motion synthesis with time-varied text prompts. MDM [65] introduces a motion diffusion model that operates on raw motion data, enabling both high-quality generation and generic conditioning that together comprise a good baseline for new motion generation tasks. Based on MDM [65], OmniControl [73] integrates flexible spatial-temporal control signals across different joints by combining analytic spatial guidance and realism guidance into the diffusion model, ensuring that the generated motion closely conforms to the input control signals. The work most relevant to ours is MLD [9], which introduces a motion latent-based diffusion model to enhance generation quality and reduce computational resource requirements. The key idea is training a VAE [30] for motion embedding, followed by implementing latent diffusion [55] within the learned latent space. However, these diffusion fashions inevitably require considerable sampling steps for motion synthesis during inference, even with some sampling acceleration methods [61]. Thus, we propose MotionLCM, which not only guarantees high-quality controllable motion generation but also achieves real-time runtime efficiency.
@@ 3.3 Controllable Motion Generation in Latent Space
$\mathbf{\Theta}^{a},\mathbf{\Theta}^{b}=\underset{\mathbf{\Theta}^{a},\mathbf{\Theta}^{b}}{\arg\min}({\cal L_{\text{recon}}}+\lambda{\cal L_{\text{{control}}}}),$ (10)
where $\lambda$ is the weight to balance the two losses. This design enables explicit control signals in the vanilla motion space to further provide supervision for the generation process. Extensive experiments demonstrate that the introduced supervision is beneficial in improving motion control performance, which will be introduced in the following section.
@@ 4.4 Ablation Studies
Table 5: Ablation study on different control ratios $\tau$ and number of control joints $K$. We report the results of (1, 2, 4)-step inference. “” is the default training setting.